OpenStudy (anonymous):

solve and check. 2x^2+3x=14

OpenStudy (wach):

We can try factoring, here, by subtracting 14 and then finding factors, right?

OpenStudy (anonymous):

yes. so it would be 2x^2+3x-14=0

OpenStudy (wach):

Correct! Now, we need to find factors of this.

OpenStudy (anonymous):

i thought it would be something like 2(x^2+?-7) but i'm not sure if that is close

OpenStudy (wach):

2x^2+3x-14 so what are 2 numbers that multiply to get 14, but when one is multiplied by 2, and then added w/ the other, we get 3?

OpenStudy (wach):

You could do it that way, but you get a fraction in the ? spot which is hard to factor

OpenStudy (anonymous):

2x7=14 would 2 and 7 work

OpenStudy (wach):

Yes! :)

OpenStudy (wach):

but one would be neg since 14 is negative

OpenStudy (wach):

now if we have to multiply one by 2 because of 2x and add it with the other one, what would the factors be? I personally think that 7 - 4 would work, yeah?

OpenStudy (wach):

(2x + 7) (x - 2) would work here since 7 x -2 = -14 , right? and since 7 + 2(-2) = 7 - 4 = 3, right?

OpenStudy (wach):

Now, set (2x + 7) = 0 and (x - 2) = 0 ans solve for x. Those are your answers.

OpenStudy (anonymous):

so the answer would be -7/2, 2 ?

OpenStudy (wach):

Yup!